OpenStudy (anonymous):

What is 3 to the power of 2 over 3 equal to? a.) cube root of 9 b.) square root of 9 c.) cube root of 27 d.) square root of 27

OpenStudy (anonymous):

I got A

OpenStudy (phi):

\[ 3^\frac{2}{3}\]

OpenStudy (phi):

you can use this idea \[ a^{bc} = \left(a^b\right)^c\] to write it as either \[ \left(3^\frac{1}{3}\right)^2 \] That means the square of the cube root of 3 or \[ \left(3^2\right)^\frac{1}{3} \] if we write 3^2 as 9 , then that is also \[ 9^\frac{1}{3} \] the cube root of 9

OpenStudy (anonymous):

@phi okay so I was right thank you!

OpenStudy (phi):

Yes, you are right. But anyone looking at this problem might want to see how to do it.

OpenStudy (anonymous):

@Phi I understand that